Under constant tension and constant mass per unit length, the note produced by a plucked string is 500 Hz when the length of the string is 0.90 m. At what length is the frequency 150 Hz?
1
Expert's answer
2020-11-10T07:03:17-0500

Frequency of the string is inversly ploportional to its length


"f =\dfrac{k}{L}"

where "k" is some constant. From the given values, let's find this constant:


"k = fL = 500Hz\times 0.9m = 450m/s"

Then for "f = 150Hz", the length will be:


"L = \dfrac{k}{f} = \dfrac{450m/s}{150Hz} = 3m"

Answer. 3m.